### v1

@@ Abstract
Anatomical trees are critical for clinical diagnosis and treatment planning, yet their complex and diverse geometry make accurate representation a significant challenge. Motivated by the latest advances in large language models, we introduce an autoregressive method for synthesizing anatomical trees. Our approach first embeds vessel structures into a learned discrete vocabulary using a VQ-VAE architecture, then models their generation autoregressively with a GPT-2 model. This method effectively captures intricate geometries and branching patterns, enabling realistic vascular tree synthesis. Comprehensive qualitative and quantitative evaluations reveal that our technique achieves high-fidelity tree reconstruction with compact discrete representations. Moreover, our B-spline representation of vessel cross-sections preserves critical morphological details that are often overlooked in previous’ methods parameterizations. To the best of our knowledge, this work is the first to generate blood vessels in an autoregressive manner. Code, data, and trained models will be made available.
@@ 1 Introduction
The choice of data representation for 3D vascular structures depends on the downstream task. However, obtaining high-fidelity reconstructions from patient scans is challenging, requiring expert knowledge and often leading to errors [13, 19]. Despite advances in vessel segmentation, accurately capturing fine vascular details remains difficult [1]. To address these limitations, various methods have been developed for synthesizing blood vessel geometries [34]. Furthermore, generative models provide a promising avenue for augmenting datasets, particularly for rare anatomical and pathological variations [4]. We identified two main approaches in the literature on generating vascular 3D models: model-based and data-driven. Model-based methods are fractal or space-filling algorithms using a set of fixed rules that include different branching parameters, generally adhering to hemodynamic laws with constraints related to flow and radius [11, 29, 25, 23]. These methods often fail to capture the complexity and diversity of real anatomical data. Instead, data-driven methods model the data distribution without the need for hard-coded rules. The first method in this line was proposed by Wolterink et al. [33], who employed Generative Adversarial Networks but restricted the model to single-channel vessels. Subsequently, Sinha et al. [27] introduced a diffusion model to learn vessel tree distributions using signed distance functions (SDFs). Deo et al. [8] extended these ideas by leveraging latent diffusion models for aneurysm segmentation generation. Kuipers et al. [12] adopted a diffusion-based approach to generate vessel point clouds but required a post-processing step to reconstruct the centerlines. Prabhakar et al. [21] proposed a novel graph-based diffusion model capable of handling cycles, though it primarily targets capillary networks and focuses on generating the vessel topology rather than a 3D mesh.
@@
VesselVAE [9], which is conceptually closer to our approach, introduced a recursive variational neural network (RvNN) tailored for vascular geometry generation. However, like other parameterized methods, it represents vessels solely by their centerlines and radii. While this compact representation is widely used, it oversimplifies the vessel’s cross-sectional geometry, potentially leading to inaccuracies in simulations based on such data [10]. Moreover, the method is limited to relatively small vascular segments, a constraint that arises from the difficulty the RvNN architecture faces when scaling to deeper levels of recursion.
@@
In light of these limitations, we propose a novel Transformer-based [32] approach for blood vessel synthesis. Although transformers have shown remarkable versatility in modeling sequential data, including geometric structures [26, 28], their application to vascular generation remains largely unexplored. In this work we present a formulation that treats vascular trees as a sequence of nodes, leveraging the GPT-2 architecture to generate compact, anatomically realistic vessel structures. Additionally, we propose using a cross-sectional representation of the vessels’ channels based on B-splines. This approach allows a more nuanced surface model compared to methods that rely solely on a single radius value, overcoming a key limitation in previous techniques. Our approach proceeds in two stages. First, we use a vector-quantized autoencoder (VQ-VAE) [31] to learn a discrete vocabulary of node attributes, capturing essential geometric features in context. We then employ this learned codebook to train a GPT-2 model [22], enabling the autoregressive generation of 3D blood vessel structures. Finally, we show extensive experiments comparing our method with baseline techniques, demonstrating superior performance and validating the effectiveness of our approach.
@@ 4 Results
We compare our method to leading vasculature generation approaches. Sinha et al. [27] uses diffusion over the space of INRs, focusing primarily on representation rather than the generation of novel structures. Feldman et al. [9] employs a recursive approach to generate vessel sequences but faces limitations with deeper trees and simplifies radius parametrization to a single value per centerline point, loosing fine details. The results are presented in Tab. 1. Both our model and the baselines were trained with two different datasets: the Aneurisk trees trimmed at height 15 and 20. For the MMD metric, while a low value indicates resemblance, an extremely small value may suggest memorization. Our model outperforms the method by [27] in both COV and 1-NNA. Although VesselVAE achieves a higher COV score, qualitative evaluation highlights its difficulties with deeper trees and its tendency to oversimplify vessel radii using circular cross-sections, preventing it from capturing fine details. While these metrics are commonly used in the literature, they fail to capture key characteristics of the data. Furthermore, Fig. 2 provides example renders for all methods and training datasets, complementing the quantitative analysis.
@@ 5 Conclusions
We introduce VesselGPT, a novel method for generating blood vessels in an autoregressive manner by sequentially sampling from a learned codebook. Our method is capable of learning a discrete optimal representation of blood vessels in an embedding space, that allows synthesizing high quality novel structures. An interesting future direction would be to evaluate the generated structures in a downstream task, such as blood flow simulations. Assessing how well the synthesized vasculature performs in practical applications would provide deeper insights into its clinical and scientific relevance. Furthermore, integrating domain-specific constraints or leveraging physics-informed models could enhance the utility and realism of the generated structures in real-world scenarios. Our approach demonstrates promising advancements in 3D blood vessel geometry synthesis, paving the way for enhanced clinical applications that could assist healthcare professionals in diagnosis, treatment planning, and surgical interventions.
@@ 5.0.1 Acknowledgements
This project was supported by grants from Salesforce, USA (Einstein AI 2020), National Scientific and Technical Research Council (CONICET), Argentina (PIP 2021-2023 GI - 11220200102981CO), and Universidad Torcuato Di Tella, Argentina.

### v2

@@ Abstract
Anatomical trees are critical for clinical diagnosis and treatment planning, yet their complex and diverse geometry make accurate representation a significant challenge. Motivated by the latest advances in large language models, we introduce an autoregressive method for synthesizing anatomical trees. Our approach first embeds vessel structures into a learned discrete vocabulary using a VQ-VAE architecture, then models their generation autoregressively with a GPT-2 model. This method effectively captures intricate geometries and branching patterns, enabling realistic vascular tree synthesis. Comprehensive qualitative and quantitative evaluations reveal that our technique achieves high-fidelity tree reconstruction with compact discrete representations. Moreover, our B-spline representation of vessel cross-sections preserves critical morphological details that are often overlooked in previous’ methods parameterizations. To the best of our knowledge, this work is the first to generate blood vessels in an autoregressive manner. Code is available at https://github.com/LIA-DiTella/VesselGPT-MICCAI.
@@ 1 Introduction
The choice of data representation for 3D vascular structures depends on the downstream task. However, obtaining high-fidelity reconstructions from patient scans is challenging, requiring expert knowledge and often leading to errors [13, 19]. Despite advances in vessel segmentation, accurately capturing fine vascular details remains difficult [1]. To address these limitations, various methods have been developed for synthesizing blood vessel geometries [34]. Furthermore, generative models provide a promising avenue for augmenting datasets, particularly for rare anatomical and pathological variations [4]. We identified two main approaches in the literature on generating vascular 3D models: model-based and data-driven. Model-based methods are fractal or space-filling algorithms using a set of fixed rules that include different branching parameters, generally adhering to hemodynamic laws with constraints related to flow and radius [11, 29, 25, 23]. These methods often fail to capture the complexity and diversity of real anatomical data. Data-driven methods model the data distribution without the need for hard-coded rules. The first method in this line was proposed by Wolterink et al. [33], who employed Generative Adversarial Networks restricted to single-channel vessels. Subsequently, Sinha et al. [27] introduced a diffusion model to learn vessel tree distributions using signed distance functions (SDFs). Deo et al. [8] extended these ideas by leveraging latent diffusion models for aneurysm segmentation generation. Kuipers et al. [12] adopted a diffusion-based approach to generate vessel point clouds but required a post-processing step to reconstruct the centerlines. Prabhakar et al. [21] proposed a novel graph-based diffusion model capable of handling cycles, though it primarily targets capillary networks and focuses on generating the vessel topology rather than a 3D mesh.
@@
VesselVAE [9], which is conceptually closer to our approach, introduced a recursive variational neural network (RvNN) tailored for vascular geometry generation. However, like other parameterized methods, it represents vessels solely by their centerlines and radii. While this compact representation is widely used, it oversimplifies the vessel’s cross-sectional geometry, potentially leading to inaccuracies in simulations based on such data [10]. Our data exploration revealed that vessel cross-sections are far from being circular, confirming the need for a non circular parametrization. Moreover, the method is limited to relatively small vascular segments, a constraint that arises from the difficulty the RvNN architecture faces when scaling to deeper levels of recursion.
@@
In light of these limitations, we propose a novel Transformer-based [32] approach for blood vessel synthesis. Although transformers have shown remarkable versatility in modeling sequential data, including geometric structures [26, 28], their application to vascular generation remains largely unexplored. In this work we present a formulation that treats vascular trees as a sequence of nodes, leveraging the GPT-2 architecture to generate compact, anatomically realistic vessel structures. Additionally, we propose using a cross-sectional representation of the vessels’ channels based on B-splines. This approach allows a more nuanced surface model compared to methods that rely solely on a single radius value, overcoming a key limitation in previous techniques. Our approach proceeds in two stages. First, we use a vector-quantized autoencoder (VQ-VAE) [31] to learn a discrete vocabulary of node attributes, capturing essential geometric features in context. We then employ this learned codebook to train a GPT-2 model [22], enabling the autoregressive generation of 3D blood vessel structures. Finally, we show experiments comparing our method with baseline techniques, validating the effectiveness of our approach.
@@ 4 Results
We compare our method to leading vasculature generation approaches. Sinha et al. [27] uses diffusion over the space of INRs, focusing primarily on representation rather than the generation of novel structures. Feldman et al. [9] employs a recursive approach to generate vessel sequences but faces limitations with deeper trees and simplifies radius parametrization to a single value per centerline point, loosing fine details. The results are presented in Tab. 1. Both our model and the baselines were trained with two different datasets: the Aneurisk trees trimmed at height 15 and 20. For the MMD metric, while a low value indicates resemblance, an extremely small value may suggest memorization. Our model outperforms the method by [27] in both COV and 1-NNA. Although VesselVAE achieves a higher COV score, qualitative evaluation highlights its difficulties with deeper trees and its tendency to oversimplify vessel radii using circular cross-sections, preventing it from capturing fine details. While these metrics are commonly used in the literature, they fail to capture key characteristics of the data. Our approach offers distinct advantages despite metric limitations: it is more efficient in training time, scales to larger and deeper vascular trees, and enables the representation of more complex anatomical features by modeling non-circular cross-sections (allowing for aneurysm modelling), an aspect not reflected by standard metrics but vital for downstream applications. Furthermore, Fig. 2 provides example renders for all methods and training datasets, complementing the quantitative analysis.
@@ 5 Conclusions
We introduce VesselGPT, a novel method for generating blood vessels in an autoregressive manner by sequentially sampling from a learned codebook. Our method is capable of learning a discrete representation of blood vessels in an embedding space, that allows synthesizing high quality novel structures. We advance the state of the art by introducing a more compact yet expressive representation of vessel geometry, using B-splines to model the cross-sectional shape. Additional metrics could be introduced in future work to better assess the impact of this representation and provide a more comprehensive evaluation of the quality of the generated structures. Assessing how well the synthesized vasculature performs in practical applications would provide deeper insights into its clinical and scientific relevance. Furthermore, integrating domain-specific constraints or leveraging physics-informed models could enhance the and realism of the generated structures in real-world scenarios. Our approach demonstrates promising advancements in 3D blood vessel geometry synthesis, paving the way for enhanced clinical applications that could assist healthcare professionals in diagnosis, treatment planning, and surgical interventions. {credits}
@@ 5.0.1 Acknowledgements
This project was supported by Universidad Torcuato Di Tella, Argentina and the National Scientific and Technical Research Council (CONICET), Argentina.